### v1

# EgoHand: Ego-centric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMUs

Abstract
Recent advanced Virtual Reality (VR) headsets, such as the Apple Vision Pro, employ bottom-facing cameras to detect hand gestures and inputs, which offers users significant convenience in VR interactions. However, these bottom-facing cameras can sometimes be inconvenient and pose a risk of unintentionally exposing sensitive information, such as private body parts or personal surroundings. To mitigate these issues, we introduce EgoHand. This system provides an alternative solution by integrating millimeter-wave radar and IMUs for hand gesture recognition, thereby offering users an additional option for gesture interaction that enhances privacy protection. To accurately recognize hand gestures, we devise a two-stage skeleton-based gesture recognition scheme. In the first stage, a novel end-to-end Transformer architecture is employed to estimate the coordinates of hand joints. Subsequently, these estimated joint coordinates are utilized for gesture recognition. Extensive experiments involving 10 subjects show that EgoHand  can detect hand gestures with 90.8% accuracy. Furthermore, EgoHand  demonstrates robust performance across a variety of cross-domain tests, including different users, dominant hands, body postures, and scenes.
1 Introduction
Figure 1: We introduce EgoHand, a VR interaction gesture recognition system utilizing head-mounted millimeter-wave radar and IMU as an alternative or complement to downward-facing cameras for enhancing personal privacy.
Virtual Reality (VR) technology has presented novel human computer interaction approaches in entertainment [28], education [3], working [5], etc. One of the most critical techniques in VR is gesture recognition. Current gesture recognition solutions mainly include joysticks [27, 1] and cameras [47, 33, 34, 45, 36]. Joysticks can provide accurate hand locomotion, while holding joysticks is an added inconvenience for the user. Cameras provide a device-free solution and have been utilized in latest VR headsets like Apple Vision Pro. However, bottom-facing cameras may expose personal privacy, such as user’s sensitive body parts, family members’ faces, and private residence environments.
To mitigate these issues, we introduce EgoHand. This system provides an alternative solution by integrating millimeter-wave radar and IMUs for hand gesture recognition, thereby offering users an additional option for gesture interaction that enhances privacy protection. Current research in mmWave radar-based gesture recognition typically involves fixing the radar at a stationary position and estimating gestures from a frontal horizontal angle [14, 25, 40, 18, 26, 37, 42, 38], Among these works, the most influential is Google’s Soli project [23, 39]. Besides, several studies have explored using mmWave radars as wearable devices to capture human movements in more dynamic settings. For example, Headar [43] integrates mmWave radar into a smartwatch to detect head nods or shakes, useful for confirmation dialogs. EgoMotion [21] and Argus [11] mount the radar on the head to perform body pose estimation and mesh reconstruction, respectively. In contrast, EgoHand  is the first to mount mmWave radar on the head for ego-centric hand pose estimation and gesture recognition.
This discrepancy renders existing mmWave radar work unsuitable for EgoHand, presenting us with two challenges. Firstly, first-person perspective VR interaction gestures encompass both single-handed and two-handed gestures, requiring automatic differentiation and recognition between them. To address this issue, we adopt a two-stage approach. The first stage utilizes end-to-end Transformer-based architectures [6, 41] to automatically estimate hand keypoint, providing end-to-end solutions capable of directly discerning single-handed or two-handed hand keypoint information. The second stage utilizes hand keypoint information as input and employs a classification approach to recognize gestures. The second challenge stems from irregular fluctuations in mmWave signals due to the movement of the mmWave device itself, as the user’s head moves. These head movements cause the radar to shift position and decrease the gesture recognition accuracy. To address this, we integrate IMUs with the mmWave radar to compensate for the positional changes caused by head motion with the feature fusion strategy.
In EgoHand, raw signals from mmWave radar are transformed into range-Doppler heatmaps and range-angle heatmaps. These, along IMUs data, are inputted separately into the mmWave Radar Encoder and IMU Encoder to learn corresponding radar representation and IMU representation, further concatenated and sequentially passed through the Pose Decoder and Context Decoder. The Pose Decoder generates a set of candidate hand pose features, while the Context Decoder further integrates temporal contextual information to produce a set of hand keypoints through MLP layers. In the training phase, we have synchronized hand movement videos and use Google MediaPipe Hand Landmark SDK to process videos into hand keypoints, later as training supervision. We apply set-based Hungarian matching algorithm [20] to optimize the MLP-generated hand keypoints with supervision. During inference, the MLP-generated hand keypoints with high confidence are fed into ResNet50 [15] to obtain the gesture recognition results in an end-to-end manner.
We conducted evaluation on 10 volunteers across 8 types of gesture interactions in 3 different scenes and postures. The mean per joint position error (MPJPE) for hand pose in EgoHand  was 72.7 mm, with a gesture recognition accuracy of 90.8%. Compared to directly utilizing ResNet50 for gesture recognition, the two-stage network approach improved recognition accuracy by approximately 30%. In cross-person experiments, the 2-shot gesture recognition accuracy was 83.9%, while 2-shot cross-scene experiments achieved 76.9% accuracy, and 2-shot cross-hand experiments reached 77.1% accuracy. These results demonstrate the effectiveness of EgoHand  in both in-domain and cross-domain VR gesture recognition. Our contributions in this work are outlined below:
(1) We introduce EgoHand, the first solution to mount mmWave radar and IMUs on the head for VR interaction gesture recognition. This approach offers advantages in user privacy protection compared to existing bottom-facing cameras, inspiring further research into privacy protection for VR hardware and applications.
(2) We propose a Transformer network for end-to-end generation of hand keypoints. Utilizing keypoints as intermediate features in a two-stage gesture recognition process, the network yields significant improvements over single-stage approaches, with accuracy enhancement up to 30%.
(3) We collect a dataset of VR interaction gestures using head-mounted mmWave radar and IMUs, totaling 26GB. The dataset has been fully open-sourced.
2 Related Work
2.1 Gesture Recognition with Cameras
In the field of computer vision, a series of methods have been developed for hand gesture recognition, most of which rely on RGB/D cameras that provide rich geometric information [31, 48]. Depth cameras perform excellently in handling complex gestures, such as the effective estimation of gesture joints through the combination of the PointNet model  [13], and Zimmermann’s proposal of the 3D Transformer network to further improve recognition accuracy  [50]. Besides, ModDrop [29] fuses video stream, depth stream, and audio stream to classify specific gestures. Zhou et al.  [49] proposed an adaptive cross-modal learning method, designing unique modal fusion strategies for different gestures to improve recognition accuracy. TMMF  [12] focuses on the single-stage recognition of continuous gestures, emphasizing the importance of temporal information in enhancing the continuity of gesture recognition.
Figure 2: EgoHand  takes the head-mounted millimeter-wave radar signals and IMU data to generate hand keypoints in the first stage, and further estimates the gesture category in the second stage. The camera is used solely for label generation during training and is not involved in inference.
2.2 Gesture Recognition with mmWave Radars
Millimeter-wave radar has significant advantages in spatial resolution and penetration ability, providing a reliable and efficient solution for human gesture recognition and pose estimation [26, 24, 30, 32]. It is noteworthy that recent research has demonstrated the progress of millimeter-wave radar in behavior detection, such as mHomeGes  [24] and mTransSee  [26], which use millimeter-wave signals to achieve real-time arm gesture recognition and environment-independent gesture recognition, respectively. mmASL  [32] extracts frequency features from 60GHz millimeter-wave signals and uses a multitask neural network to recognize american sign language. Pantomime et.al. [30] uses millimeter-wave radar to compute sparse 3D point clouds for gesture recognition in their self-collected dataset.
However, most of these mmWave radar based systems are designed to be fixed installations; adapting them to a user’s free movement in space, and overcoming limitations of viewpoint and location remain challenges, which may limit their application in VR interactions.
3 Methods
3.1 Data Preprocessing
In our setup, mmWave radar transmits and receives frequency modulated continuous wave (FMCW) signals chirp by chirp (64 chirps per frame, 20 frames per seconds), and mixes the received signals with the transmitted signals to obtain an intermediate frequency (IF) signal [2, 17, 9]. Due to the sparse point cloud in mmWave radar data, typically consisting of only a few points, hand reflections are inadequately represented, and minor posture changes can cause significant variations, leading to inconsistent representations. To address this issue, we adopted a preprocessing method similar to that in mHomeGes [24] to obtain a richer representation of the radar data.
$\bullet$ Range FFT. We apply range FFT on every IF chirp, which reveals the frequency difference between the received and transmitted signals, to obtain distance heatmaps of reflection objects.
$\bullet$ Doppler FFT. There exists a phase difference between neighboring chirps caused by the object movements, we use Doppler FFT on chirps in one frame along the direction of the phase change caused by the Doppler effect to estimate the object’s velocity, the outputs called range-Doppler heatmaps.
$\bullet$ Angle FFT. There are phase differences between the antennas caused by the spatial location. We use angle FFT is executed on the range-Doppler heatmaps along the receiving antenna dimension to acquire range-angle heatmaps, which responds to the distance and angle information of the reflection objects.
We concatenate range-Doppler heatmaps and range-angle heatmaps along the vertical axis for the deep network input, as shown in Fig. 3. In contrast, we do not preprocess IMU series.
Figure 3: Millimeter-wave radar signal processing involves several Fourier transform steps focusing on the time domain, chirp signals, and the receiving antenna dimension.
3.2 Deep Network Design
We use an 15.85 mW low-power mmWave radar to capture hand movements. However, the radar signal also contains noise from head movements. To compensate for this, we attach an IMU to the mmWave radar to record its motion. This creates a multimodal fusion problem, and recent research shows that Transformer architectures perform well in multimodal learning [7, 22]. Therefore, we use a Transformer-based structure for representation learning.
Another design choice arose from experimental results: directly predicting gesture classes led to poor performance, with accuracy around 65%. We therefore opted to first generate an intermediate representation as a transition step, then use it to classify gestures. Research using cameras [8, 16] has shown that hand skeletons serve as a promising and interpretable intermediate representation, so we adopted this approach.
The third design consideration is handling gestures involving different numbers of hands: 0, 1, or 2 hands may be present during interactions. Unlike OCHID-Fi [46], which assumes a fixed output of keypoints for one hand, we require an approach that can adapt to varying numbers of hands. This is similar to the object detection task, where the model must flexibly output coordinates for a variable number of targets. Therefore, we adapt DETR [6], an end-to-end object detection framework, to regress the coordinates of hand keypoints.
Based on the above considerations, we propose the framework shown in Fig. 2 and the network structure illustrated in Fig. 4. More detailed operations will be introduced next.
Figure 4: mmVR deep network consists of four main novel components. (a) mmWave Radar Encoder, (b) IMU Encoder, (c) Pose Decoder, and (d) Context Decoder.
$\bullet$ mmWave Radar Encoder. This encoder integrates six encoder blocks to process mmWave radar streams for mmWave radar embeddings. Each block comprises a multi-head self-attention module and a Feed-Forward Network (FFN), which are fundamental components of the Transformer architecture  [35].
$\bullet$ IMU Encoder. This part of the architecture is similar to the mmWave Radar Encoder, which is used for IMU data embeddings. The encoded IMU embeddings are concatenated with the mmWave radar embeddings for further Pose Decoder and Context Decoder.
$\bullet$ Pose Decoder. The structure of the decoder layer is the basic module of the DETR [6] decoder, and we stacked three such layers. One Pose Decoder block contains one multi-head self-attention module and one multi-head cross-attention module. The multi-head cross-attention module compute attention matrix between Pose Queries ($Q_{pose}=[q_{1},q_{2},\ldots,q_{C}]\in\mathbb{R}^{C\times D}$) and the fused embeddings ($E_{fused}$) from mmWave Radar Encoder and IMU Encoder. Pose Decoder outputs $X_{pose}\in\mathbb{R}^{C\times D}$, representing features of $C$ hand candidates in $D$ dimension. The process of Pose Decoder can be formulated as follows:
$X_{pose}=\text{PoseDecoder}(Q_{pose},E_{fused})$ (1)
$\bullet$ Context Decoder. One Context Decoder block includes one multi-head self-attention module and two multi-head cross-attention modules, and we stack 30 such blocks in the Context Decoder. Context Decoder learns the globally temporal relationships among different frames to refine the hand keypoint estimation. Specifically, Context Decoder utilizes the multi-head cross-attention to compute cross-attention matrix between the current query and previous pose features from Pose Decoder, iteratively. At last, two Multi-Layer Perceptrons (MLPs) are respectively to generate the keypoints of the hand candidates and the corresponding confidence scores indicating the authenticity of these hand candidates. The process of Context Decoder can be formulated as follows:
$H_{set},H_{score}=\text{ContextDecoder}(Q_{context},E_{fused},X_{pose})$ (2)
where $Q_{context}\in\mathbb{R}^{C\times D}$ represents $C$ learnable context queries; $H_{set}\in\mathbb{R}^{C\times n\times K\times 3}$ represents K 3D-keypoints of C hand candidates of $n$ frames. $H_{score}\in\mathbb{R}^{C\times n}$ represents confidence scores of these hand candidates in $n$ frames. In our settings, $C=100$, $K=21$, and $n=30$.
$\bullet$ Loss Function. We implement a Hungarian Matching algorithm [20] to ensure a unique prediction for the ground-truth 3D hand keypoints of each hand candidate.
$L_{kpt}=\text{HungarianMatch}(H_{set},H_{gt})$ (3)
where $H_{gt}$ denotes the ground-truth hand keypoint coordinates. $H_{set}$ is selected up to two hand candidates based on $H_{score}$. If there are two hands, their keypoints are concatenated and fed into the recognition network. If there is only one hand, it is duplicated before concatenation, ensuring that the recognition network has a fixed input size.
Further, we apply a cross-entropy function to compute the gesture recognition loss.
$L_{cls}=\text{CrossEntropy}(y,\hat{y})$ (4)
where $y$ and $\hat{y}$ represent the gesture label and recognition confidence score, respectively.
The total loss to optimize the deep network is a weighted combination of hand keypoint regression loss and the gesture regression loss.
$Loss=L_{kpt}+\lambda L_{cls}$ (5)
where $\lambda$ is a weight parameter balancing the hand keypoint regression and gesture recognition.
$\bullet$ Training Details. We employed the Adam optimizer [19] to train deep network on a NVIDIA 3090 GPU. The batch size was 32. The training was carried out 500 epochs. The initial learning rate was 0.001 and was halved every 100 epochs. At the start of each epoch, the training samples were randomly shuffled. To balance the keypoint regression and gesture recognition, we experimentally set $\lambda=1$ in Equation. 5.
4 Experiments
4.1 Dataset Acquisition
Dataset was collected under the approval of IRB.
(1) Hardware Configurations. The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, shown in Fig. 5. These devices were manually synchronized before recording started.
$\bullet$ Millimeter-wave radar. One TI IWR6843ISK radar generates frequency modulated continuous wave (FMCW) signals at frequencies of 60GHz-64GHz. Its transmit power is only 15.85 mW, slightly lower than typical Wi-Fi devices, making it safe for human exposure. The radar has three transmitting antennas and four receiving antennas. The transmit parameters are set to 20 frames per second with 64 chirps per frame. The analog-to-digital converter (ADC) sampling rate was 256, and the raw radar data was obtained by mounting the radar on a TI DCA1000EVM board. The raw radar data is $\in 20t\times 3\times 4\times 256\times 64$, where $t$ is the time in seconds. The range-Doppler and range-angle heatmaps are obtained after the preprocessing process, both of which are $\in 20t\times 256\times 64$.
$\bullet$ IMU. The IMU sensor records the acceleration and angular velocity data in the X, Y, and Z axes at a rate of 20 samples per second. The recorded samples are $\in 20t\times 2\times 3$, where $t$ is the time in seconds.
$\bullet$ Camera. A Hikvision camera captures the RGB videos at 20 frames per second in $1080\times 720$ resolution.
(2) Dataset diversity. We recruited 10 volunteers with heights ranging from 161 to 182 centimeters and weights between 44 to 73 kilograms. We conducted our experiment across three different scenes and three distinct postures, i.e., sitting, lying, and standing, wherein each volunteer engaged in 8 types of human-computer interaction gestures. These gestures include click, swipe leftward, swipe rightward, swipe upward, swipe downward, swipe to bottom right, zoom in, and zoom out. The latter two are two-handed gestures and the rest are single-handed gestures. The selection of these gestures and postures is inspired by the Apple Vision Pro, such as confirming selection and scrolling up the page, which has demonstrated that this set of gestures can facilitate a wide range of VR interactive functions, as shown in the promotional videos released by Apple [^1] Youtube link and Bilibili link. In each scene and posture, volunteers were instructed to perform each gesture 20 times, with each gesture completed within 2 seconds.
Of the 10 volunteers, two are left-handed, and in performing single-handed gestures, they have been assigned tasks to execute using their left and right hands, respectively. The remaining eight volunteers are right-handed, among whom six have been arranged to carry out interactive gestures using their dominant right hand, while two have been assigned to perform gestures using their non-dominant left hand. Thus, there are $2\times 2+6+2=12$ different settings for the left and right hand executing gestures in total. Table. 1 shows this experimental setting. This setup has been designed for scenarios where a user’s preferred hand is occupied, ensuring that the system can still recognize and respond to gestures executed by the non-preferred hand.
No. ID Pre.Hand Exe.Hand Gender H.(cm) W.(kg) BMI
1 1 Left Right Female 173 58 19.38
2 1 Left Left Female 173 58 19.38
3 2 Left Right Male 175 73 23.84
4 2 Left Left Male 175 73 23.84
5 3 Right Left Male 172 62 20.96
6 4 Right Right Female 165 45 16.53
7 5 Right Right Male 182 67 20.23
8 6 Right Right Male 174 68 22.46
9 7 Right Right Male 161 44 16.97
10 8 Right Right Male 178 65 20.52
11 9 Right Right Male 175 50 16.33
12 10 Right Left Female 163 48 18.07
Table 1: We recruited 10 volunteers to perform VR interaction gestures. They were asked to execute gestures by their preferred hand or non-preferred hand. Pre.Hand for Preferred hand; Exe.Hand for Executing Hand; H. for Height; W. for Weight.
Figure 5: Hardware setup. The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, including three postures in three scenes, i.e., standing at scene #3, lying at scene #2, and sitting at scene #1, respectively.
(3) Dataset statistics. The dataset contains 5760 (3 scenes $\times$ 12 hand execuring settings $\times$ 8 gestures $\times$ 20 repetitions) gesture samples, including synchronized data of millimeter-wave radar, IMUs, and videos. We uniformly select the 5th, 10th, 15th, and 20th repetitions for test, and the remaining 16 repetitions for training. We utilize Google MediaPipe Hand Landmark SDK to extract 3D hand keypoints from videos as the ground truth for deep network training and evaluation. We observe that the SDK sometimes misses to extract hand keypoints, especially in situations when hands move very fast. If the SDK misses 40% or more frames in a particular video, we discard it. Otherwise, we downsample or upsample the hand keypoints to 30 frames per gesture sample. Although MediaPipe’s 3D positions may lack high precision, this approach significantly simplifies the annotation process. Consequently, we have a dataset that consists of 5206 gesture samples, detailed statistics shown in Table. 2.
sitting in s#1 lying in s#2 standing in s#3 all
#training 1384 1384 1384 4152
#test 350 351 353 1054
Table 2: Dataset samples. Volunteers executed the gestures with postures of sitting in scene #1, lying in in scene #2, and standing at scene #3, respectively.
4.2 Evaluation Metrics
(1) Mean Per Joint Position Error(MPJPE). The L2 norm is computed between the predicted and true 3D coordinates of the hand joints, as shown in Equation 6. This computation leads to the Joint Position Error.
$\operatorname{PJPE}(\mathrm{k})=\frac{1}{F}\sum_{f=1}^{F}|pre(f,k)-gt(f,k)|_{2}$ (6)
$\operatorname{MPJPE}=\frac{1}{K}\sum_{k=1}^{K}E_{PJPE}(k)$ (7)
$\operatorname{PJPE}(\mathrm{k})$ is the PJPE for the k-th joint. The MPJPE is the average of all $\operatorname{PJPE}(\mathrm{k})$.
(2) We also adopt gesture recognition Accuracy, Precision, Recall, and F1-score for the evaluation.
4.3 In-domain Results
Overall performance. Table. 3 presents the main results of EgoHand  under the data splitting strategy shown in Table. 2. The one-stage approaches directly output the gesture classification results without regressing the hand keypoints. In contrast, EgoHand  outputs hand keypoints as the intermediate features and conduct gesture recognition over the keypoints. We evaluate current mainstream and basic classification networks. Table. 3 demonstrates that EgoHand  outperforms one-stage approaches by a large margin (except EgoHand  (ViT [10])). This result demonstrates that the novel end-to-end Transformer architecture proposed in our first stage is highly effective. A more intriguing finding is that, despite not achieving particularly good MPJPE in the first stage, intermediate hand keypoint positions significantly boost the accuracy of gesture recognition. This suggests that future work could consider promoting hand keypoint estimation to enhance gesture recognition. In light of the good performance of ResNet50 in classification tasks, this network has been selected as the gesture recognition backbone.
Metrics MPJPE Acc(%) Precision Recall F1
one-stage (LSTM) / 67.99 70.58 68.35 67.80
one-stage (ViT) / 59.66 67.12 59.77 60.75
one-stage (ResNet50) / 60.09 61.79 61.19 60.64
one-stage (mGesNet) / 81.34 81.98 81.56 81.64
one-stage (mSeeNet) / 84.60 86.73 86.59 86.60
one-stage (mmGesture) / 68.66 70.01 68.83 68.81
EgoHand (GCN) 81.95 85.36 87.16 87.12 87.09
EgoHand (LSTM) 75.70 89.26 91.56 91.48 91.46
EgoHand (ViT) 91.20 77.95 81.39 80.19 79.28
EgoHand (ResNet50) 72.73 90.80 93.34 93.20 93.19
Table 3: EgoHand outperforms one-stage methods such as mGesNet [24], mSeeNet [26], and mmGesture [40], by a large margin. MPJPE is measured in millimeters(mm).
Gesture Acc(%) Precision Recall F1
Click 99.21 96.15 99.21 97.66
Swipe leftward 97.78 96.35 97.78 97.06
Swipe rightward 97.06 96.35 97.06 96.70
Swipe upward 88.43 97.27 88.43 92.64
Swipe downward 94.74 95.45 94.74 95.09
Swipe to bottom right 100.00 97.10 100.00 98.53
Zoom in 79.51 87.39 79.51 83.26
Zoom out 88.89 80.62 88.89 84.55
mean 90.80 93.34 93.20 93.19
Table 4: Accuracy, Precision, Recall and F1 for each gesture.
Figure 6: EgoHand  hand pose estimation examples, showing click, swipe downward, zoom in, and swipe leftward, respectively. (Images here are only used for visualization).
PJPEs for fingers. In addition, PJPEs of the thumb finger, index finger, middle finger, ring finger, and little finger are 60.18, 70.10, 76.42, 84.44, and 91.79, respectively. The thumb and index finger exhibit the least prediction errors, whereas the largest errors are found in the ring and little fingers. This is because, from the perspective of a mmWave radar worn on the head, the thumb and index finger are the most discernible, while the other fingers tend to be frequently obscured by the thumb and index finger.
Visualization. Fig. 6 illustrates the visualization of hand keypoint predictions by the EgoHand  for various gestures, showcasing different volunteers performing gestures such as click, swipe downward, zoom in, and swipe leftward. Each set of gestures is depicted in three rows: the first row presents the video frames captured during the gestures, the second row shows the hand keypoints derived from the video frames using the Google SDK, and the third row displays the predicted hand keypoints output by EgoHand. It can be seen that EgoHand  can automatically detect the number of hands, whether they are left or right hands, as well as the movement of hand joints throughout the gesture executing process, which effectively alleviates the difficulty of subsequent gesture recognition.
Gesture recognition. Table. 4 shows the Accuracy, Precision, Recall and F1 Score for each gesture. We can see that the accuracy of single-hand gestures is generally above 90%, whereas the accuracy of two-hand gestures is slightly lower but still remains above 80%. The most dominant classification errors are the mutual misclassifications between the two-hand gestures of zoom in and zoom out. This is because, during the data collection, volunteers have no strict guidance on which diagonal direction should be used for zoom in and zoom out, introducing more variability compared to single-hand gestures. Besides, The MPJPEs of zoom in and zoom out in the first stage are generally higher than the single-hand gestures, leading to suboptimal performance in the subsequent gesture classification.
Modality ablation study. Table. 5 shows the experiments relating to the impact of IMU data in EgoHand. The experiments employed strategies using only millimeter-wave radar data, only IMU data, and both data streams. Results indicate that using only IMU data to estimate gestures results in weakest performance in terms of both mean MPJPE and gesture recognition accuracy. Relying only on mmWave radar data also leads to a decrease in metrics compared to the both using millimeter-wave radar and IMU data. This experiment demonstrates that data captured by IMUs can effectively compensate for human body or head movements, thereby enhancing the accuracy of hand keypoint estimation as well as gesture recognition.
4.4 Cross-domain Evaluation
Cross-person. We utilize two cross-person validation approaches to evaluate the performance of EgoHand  in gesture recognition on unseen users. (1) We use dataset from No.1 to No.8 listed in Table. 1 as the training set to train EgoHand, and test the trained EgoHand  with dataset from No.9 to No.12. (2) We select one of 12 rows from Table. 1 for testing and use the remaining rows for training; and we iteratively repeat this process 12 times and ultimately compute the average value. Table. 6 shows that EgoHand  maintains a certain level of recognition capability for unseen persons, achieving a gesture recognition accuracy of 63.11 and 61.96% in the above two training/test splitting strategy, respectively. Moreover, the gesture recognition accuracy significantly improves by 10%-20% with one-shot or two-shot fine-tuning, using only one or two samples from the target person/s to fine-tune EgoHand.
Input MPJPE Acc(%) Precision Recall F1
IMUs 120.18 68.80 70.13 69.80 69.36
mmWave 96.42 86.04 88.70 88.55 88.54
Fusion 72.73 90.80 93.34 93.20 93.19
Table 5: With mmWave radar and IMUs, EgoHand  performs the best. MPJPE is measured in millimeters(mm).
#Training; #Test #-Shot Acc(%) Precision Recall F1
zero 63.11 64.42 63.17 62.74
No.1-8; No.9-12 (Table. 1) one 73.99 74.50 73.60 73.57
two 83.97 84.13 83.88 83.87
zero 61.96 68.16 65.54 64.68
leave-one-person-out one 69.37 74.97 74.41 73.85
two 74.65 77.86 76.51 75.09
Table 6: Gesture recognition tests on unseen volunteers. EgoHand  maintains a certain level of recognition capability for unseern persons, and can be improved by 10%-20% with one-shot or two-shot fine-tuning.
Figure 7: Gesture recognition accuracy under leave-one-person-out setting.
#Training; #Test #-Shot Acc(%)
sitting in scene #1 (in-domain result dataset split) / 87.60
lying in scene #2 (in-domain result dataset split) / 87.96
standing in scene #3 (in-domain result dataset split) / 95.92
zero 54.35
scene #2 (lying) and #3 (standing); scene #1 (sitting) one 67.58
two 74.97
zero 51.42
scene #1 (sitting) and #3 (standing); scene #2 (lying) one 70.33
two 76.97
zero 51.41
scene #1 (sitting) and #2 (lying); scene #3 (standing) one 60.42
two 74.51
Table 7: Gesture recognition tests on unseen posture/scene. EgoHand  maintains a certain level of recognition capability for unseen posture/scene, and can be improved by 20% with one-shot or two-shot fine-tuning.
Fig. 7 demonstrates the leave-one-person-out results. It also shows that EgoHand  maintains a certain level of recognition capability and can be improved further by few-shot fine-tuning. There is one exception observed with No. 7, which exhibits a significantly lower accuracy rate. Upon reviewing the video footage, we find that this particular volunteer performs the gestures with smaller and more casual movements, deviating considerably from the patterns exhibited by others. Consequently, during actual system deployment, what we need is to implement a pre-use calibration protocol, requiring users to calibrate each gesture twice or more before commencing usage.
Cross-posture and Cross-scene. We use data from two postures (or scene) in Table. 2 to train EgoHand and tests the trained EgoHand on data of the last postures (or scene). Similar to the cross-person performance, EgoHand  maintains a certain level of recognition capability for unseen postures/scenes, achieving a gesture recognition accuracy of 54.35%, 51.42%, and 51.41% in three cross-posture evaluation, respectively. Moreover, the gesture recognition accuracy significantly improves by 20% with one-shot or two-shot fine-tuning. This experiment further substantiates the efficiency of pre-use calibration strategy discussed in the above cross-person evaluation.
Data splitting strategy #Training Hand #Test Hand #-Shot Acc(%)
In-domain results accuracy of left hand / 90.96
accuracy of right hand / 93.04
Cross-hand evaluation zero 28.25
Right Left one 57.80
two 73.53
zero 24.95
Left Right one 58.92
two 77.18
Table 8: Gesture recognition tests on unseen hands. EgoHand  maintains a weak level of recognition capability for unseern hands, and can be largely improved by 30%-50% with one-shot or two-shot fine-tuning.
Cross-hand. Table. 8 shows that EgoHand  achieves that the accuracy for each hand exceeds 90% if following the data splitting strategy in in-domain results section. We also conduct cross-hand evaluation, training EgoHand  with data of right/left executing hand in Table. 1, and test the trained EgoHand  with data of left/right executing hand, respectively. It is noteworthy that this training strategy also encompasses cross-person challenges. For example, as shown in Table. 1, instances from No.6 to No.11 introduce cross-person challenges when the left-hand execution data is used for training and the right-hand execution data is employed for testing. Thus, this is a highly challenging data splitting scheme, with Table. 8 indicating that EgoHand  struggles greatly under such split. However, upon fine-tuning with just one to two samples, gesture recognition accuracy sees a remarkable improvement of 30-50%, up to 73.53% and 77.18%. This indicates that during the training of EgoHand, we should simultaneously provide samples from both the left and right hands.
5 Conclusion
In this paper, we present EgoHand, an proof-of-concept solution for VR interaction gesture recognition utilizing millimeter-wave radar and IMUs worn on the head for preserving more personal privacy. We conduct extensive cross-domain and few-shot experiments to evaluate EgoHand, highlighting the importance of training with samples from both hands to enhance cross-hand stability. Additionally, at least two pre-use calibrations are essential to ensure optimal performance across various users, scenes, and postures.
Current EgoHand appears somewhat cumbersome, particularly with the TI mmWave radar. This can potentially be improved, for example, by using Google’s Soli radar (9mm×$\times$9mm) [23]. In addition, mmWave radar may expose health metrics like breathing and heart rate [44, 4], which can be mitigated by adjusting the range and angle of radar signals.

### v2

@@
-EgoHand: Ego-centric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMUs
+mmEgoHand: Egocentric Hand Pose Estimation and Gesture Recognition with Head-mounted Millimeter-wave Radar and IMU
 Abstract
-Recent advanced Virtual Reality (VR) headsets, such as the Apple Vision Pro, employ bottom-facing cameras to detect hand gestures and inputs, which offers users significant convenience in VR interactions.
-However, these bottom-facing cameras can sometimes be inconvenient and pose a risk of unintentionally exposing sensitive information, such as private body parts or personal surroundings.
-To mitigate these issues, we introduce EgoHand.
-This system provides an alternative solution by integrating millimeter-wave radar and IMUs for hand gesture recognition, thereby offering users an additional option for gesture interaction that enhances privacy protection.
-To accurately recognize hand gestures, we devise a two-stage skeleton-based gesture recognition scheme.
-In the first stage, a novel end-to-end Transformer architecture is employed to estimate the coordinates of hand joints.
-Subsequently, these estimated joint coordinates are utilized for gesture recognition.
-Extensive experiments involving 10 subjects show that EgoHand  can detect hand gestures with 90.8% accuracy.
-Furthermore, EgoHand  demonstrates robust performance across a variety of cross-domain tests, including different users, dominant hands, body postures, and scenes.
+Recent advancements in millimeter-wave (mmWave) radar have demonstrated its potential for human action recognition and pose estimation, offering privacy-preserving advantages over conventional cameras while maintaining occlusion robustness, with promising applications in human-computer interaction and wellness care.
+However, existing mmWave systems typically employ fixed-position configurations, restricting user mobility to predefined zones and limiting practical deployment scenarios.
+We introduce mmEgoHand, a head-mounted egocentric system for hand pose estimation to support applications such as gesture recognition, VR interaction, skill digitization and assessment, and robotic teleoperation.
+mmEgoHand  synergistically integrates mmWave radar with inertial measurement units (IMUs) to enable dynamic perception.
+The IMUs actively compensate for radar interference induced by head movements, while our novel end-to-end Transformer architecture simultaneously estimates 3D hand keypoint coordinates through multi-modal sensor fusion.
+This dual-modality framework achieves spatial-temporal alignment of mmWave heatmaps with IMUs, overcoming viewpoint instability inherent in egocentric sensing scenarios.
+We further demonstrate that intermediate hand pose representations substantially improve performance in downstream task, e.g., VR gesture recognition.
+Extensive evaluations with 10 subjects performing 8 gestures across 3 distinct postures- standing, sitting, lying - achieve 90.8% recognition accuracy, outperforming state-of-the-art solutions by a large margin.
+Dataset and code are available at https://github.com/WhisperYi/mmVR.
+Index Terms:
+Human sensing, Hand pose estimation, Gesture recognition, Millimeter-wave radar, IMUs, Human-computer interaction
 1 Introduction
-Figure 1: We introduce EgoHand, a VR interaction gesture recognition system utilizing head-mounted millimeter-wave radar and IMU as an alternative or complement to downward-facing cameras for enhancing personal privacy.
-Virtual Reality (VR) technology has presented novel human computer interaction approaches in entertainment [28], education [3], working [5], etc.
-One of the most critical techniques in VR is gesture recognition.
-Current gesture recognition solutions mainly include joysticks [27, 1] and cameras [47, 33, 34, 45, 36].
-Joysticks can provide accurate hand locomotion, while holding joysticks is an added inconvenience for the user.
-Cameras provide a device-free solution and have been utilized in latest VR headsets like Apple Vision Pro.
-However, bottom-facing cameras may expose personal privacy, such as user’s sensitive body parts, family members’ faces, and private residence environments.
-To mitigate these issues, we introduce EgoHand.
-This system provides an alternative solution by integrating millimeter-wave radar and IMUs for hand gesture recognition, thereby offering users an additional option for gesture interaction that enhances privacy protection.
-Current research in mmWave radar-based gesture recognition typically involves fixing the radar at a stationary position and estimating gestures from a frontal horizontal angle [14, 25, 40, 18, 26, 37, 42, 38], Among these works, the most influential is Google’s Soli project [23, 39].
-Besides, several studies have explored using mmWave radars as wearable devices to capture human movements in more dynamic settings.
-For example, Headar [43] integrates mmWave radar into a smartwatch to detect head nods or shakes, useful for confirmation dialogs.
-EgoMotion [21] and Argus [11] mount the radar on the head to perform body pose estimation and mesh reconstruction, respectively.
-In contrast, EgoHand  is the first to mount mmWave radar on the head for ego-centric hand pose estimation and gesture recognition.
-This discrepancy renders existing mmWave radar work unsuitable for EgoHand, presenting us with two challenges.
-Firstly, first-person perspective VR interaction gestures encompass both single-handed and two-handed gestures, requiring automatic differentiation and recognition between them.
-To address this issue, we adopt a two-stage approach.
-The first stage utilizes end-to-end Transformer-based architectures [6, 41] to automatically estimate hand keypoint, providing end-to-end solutions capable of directly discerning single-handed or two-handed hand keypoint information.
-The second stage utilizes hand keypoint information as input and employs a classification approach to recognize gestures.
-The second challenge stems from irregular fluctuations in mmWave signals due to the movement of the mmWave device itself, as the user’s head moves.
-These head movements cause the radar to shift position and decrease the gesture recognition accuracy.
-To address this, we integrate IMUs with the mmWave radar to compensate for the positional changes caused by head motion with the feature fusion strategy.
-In EgoHand, raw signals from mmWave radar are transformed into range-Doppler heatmaps and range-angle heatmaps.
-These, along IMUs data, are inputted separately into the mmWave Radar Encoder and IMU Encoder to learn corresponding radar representation and IMU representation, further concatenated and sequentially passed through the Pose Decoder and Context Decoder.
-The Pose Decoder generates a set of candidate hand pose features, while the Context Decoder further integrates temporal contextual information to produce a set of hand keypoints through MLP layers.
-In the training phase, we have synchronized hand movement videos and use Google MediaPipe Hand Landmark SDK to process videos into hand keypoints, later as training supervision.
-We apply set-based Hungarian matching algorithm [20] to optimize the MLP-generated hand keypoints with supervision.
-During inference, the MLP-generated hand keypoints with high confidence are fed into ResNet50 [15] to obtain the gesture recognition results in an end-to-end manner.
-We conducted evaluation on 10 volunteers across 8 types of gesture interactions in 3 different scenes and postures.
-The mean per joint position error (MPJPE) for hand pose in EgoHand  was 72.7 mm, with a gesture recognition accuracy of 90.8%.
-Compared to directly utilizing ResNet50 for gesture recognition, the two-stage network approach improved recognition accuracy by approximately 30%.
-In cross-person experiments, the 2-shot gesture recognition accuracy was 83.9%, while 2-shot cross-scene experiments achieved 76.9% accuracy, and 2-shot cross-hand experiments reached 77.1% accuracy.
-These results demonstrate the effectiveness of EgoHand  in both in-domain and cross-domain VR gesture recognition.
-Our contributions in this work are outlined below:
-(1) We introduce EgoHand, the first solution to mount mmWave radar and IMUs on the head for VR interaction gesture recognition.
-This approach offers advantages in user privacy protection compared to existing bottom-facing cameras, inspiring further research into privacy protection for VR hardware and applications.
-(2) We propose a Transformer network for end-to-end generation of hand keypoints.
-Utilizing keypoints as intermediate features in a two-stage gesture recognition process, the network yields significant improvements over single-stage approaches, with accuracy enhancement up to 30%.
-(3) We collect a dataset of VR interaction gestures using head-mounted mmWave radar and IMUs, totaling 26GB. The dataset has been fully open-sourced.
+Figure 1: We present mmEgoHand, an interaction gesture recognition system utilizing head-mounted millimeter-wave radar and IMUs as an alternative or complement to downward-facing cameras for enhancing personal privacy.
+Human activity recognition and pose estimation are foundational technologies that support a wide range of applications, including human behavior analysis, wellness care, human-computer interaction, and gaming.
+Traditionally, these tasks have relied on camera-based solutions [1, 2].
+More recently, wireless sensing alternatives such as Wi-Fi [3] and custom-built radar systems [4] have emerged to address the limitations of visual methods, offering improved privacy and robustness to occlusion.
+However, Wi-Fi-based approaches are sensitive to environmental variations and subject orientation [5, 6], while custom radar systems face challenges in adoption due to hardware accessibility.
+These constraints have spurred interest in commercially available millimeter-wave (mmWave) radar systems, such as the Texas Instruments (TI) IWR/AWR series, which offer a compelling trade-off between deployment practicality and high-resolution motion sensing for human activity recognition and pose estimation.
+The evolution of mmWave radar-based human sensing has progressed from coarse activity recognition [7, 8, 9, 10, 11, 12, 13] to fine-grained pose estimation [14, 15, 16, 17, 18, 19, 20, 21, 22].
+Most existing systems employ frontal, static deployments, requiring users to stay within a constrained sensing volume—an approach that limits applicability in dynamic environments.
+Recent advances in egocentric sensing, such as mmEgo [23] and Argus [24], have begun to overcome this limitation by leveraging head-mounted radar to capture body-reflected signals for pose estimation.
+These systems enable full-body tracking while preserving human movement, shifting mmWave sensing from environment-anchored perception to mobile, user-centric motion capture.
+In this work, we introduce mmEgoHand, an egocentric hand pose estimation system.
+While prior mmWave methods primarily focus on full-body tracking, our system targets fine-grained hand articulation to support applications such as gesture recognition, VR interaction, skill digitization and assessment, and robotic teleoperation.
+The design of mmEgoHand  is guided by three key considerations: (1) Existing frontal-view systems [25, 26, 27] typically estimate single-hand poses, whereas many real-world gestures involve both hands.
+We thus design a unified framework capable of handling both single- and two-hand configurations.
+(2) Head-mounted radar introduces ego-motion artifacts, as head movements dynamically alter the radar’s perspective, changing signal characteristics for the same hand gesture.
+Our system incorporates active motion compensation to disentangle hand motion from head-induced signal variations.
+(3) Traditional motion capture systems, such as VICON [28], are cumbersome for annotating hand poses.
+We leverage a lightweight annotation pipeline that ensures high-quality ground truth while significantly simplifying deployment, improving adaptability across application scenarios.
+To address the first consideration, we adapt an end-to-end Wi-Fi-based multi-person pose estimation [29].
+In the original design, the system automatically detects a variable number of individuals and estimates their body keypoints.
+Similarly, mmEgoHand  is designed to output a variable number of hands with a set-based Hungarian matching algorithm [30] and estimate keypoints for each detected hand in an end-to-end manner.
+This modification enables mmEgoHand  to support both one-handed and two-handed interaction tasks, making it more versatile and applicable to a wider range of scenarios compared to existing methods such as mm4Arm [27] and mmHand [26, 25], which only support fixed one-handed interactions.
+Furthermore, we augment the original architecture with a context decoder that takes a sequence of mmWave heatmaps as input, leveraging temporal information to improve the accuracy of hand pose estimation.
+For the second consideration, we draw inspiration from the hardware setup of mmEgo [23], attaching an inertial measurement unit (IMU) near the mmWave radar to capture head motion.
+The IMU data is temporally synchronized with the mmWave input and fed jointly into the network.
+By fusing these two modalities, mmEgoHand  can compensate for ego-motion artifacts caused by head movement, leading to more stable and accurate hand pose estimation.
+To address the third consideration, we employ a lightweight annotation strategy during data collection.
+A standard web camera is placed next to the subject to record hand movements, and hand keypoints are automatically extracted from the video using Google MediaPipe Hand Landmark SDK.
+We then perform extensive manual checking and filtering, removing keypoints with poor tracking quality, particularly under fast hand motion conditions, and take the remaining as ground truth annotations for training and evaluation.
+Similar camera-based automatic annotation schemes have become increasingly popular in prior work [3, 4, 24], offering a practical trade-off between annotation quality and deployment simplicity.
+To evaluate mmEgoHand, we recruited 10 volunteers who performed hand movements in three different scenes and under three distinct postures: standing, sitting, and lying.
+mmEgoHand  achieved a mean per-joint position error (MPJPE) of 72.73mm for hand pose estimation.
+In comparison, using mmWave radar alone resulted in an MPJPE of 96.42mm, highlighting the effectiveness of incorporating IMU data.
+Furthermore, we used the estimated hand poses from mmEgoHand  as intermediate representations for a downstream hand gesture recognition task.
+mmEgoHand  achieved an accuracy of 90.80%, substantially outperforming state-of-the-art approaches such as mGesNet [7] (81.34%), mSeeNet [9] (84.60%), and mmGesture [31] (68.66%).
+These results highlight that mmEgoHand  not only enhances hand pose estimation, but also provides more reliable and informative features for downstream applications.
+Our contributions are summarized as follows:
+(1) We introduce mmEgoHand, the first egocentric system capable of human hand pose estimation and gesture recognition (see Table I for a comparison with prior work).
+(2) We propose a novel dual-decoder architecture that simultaneously addresses spatial information and temporal coherence, yielding significant improvements over state-of-the-art approaches
+(3) We collect and open-source a large-scale dataset of VR interaction gestures captured using head-mounted mmWave radar and IMU, totaling 26GB, to support future research in human sensing.
+Dataset and code are available at https://github.com/WhisperYi/mmVR
+TABLE I: mmEgoHand  is the first egocentric system capable of simultaneously capturing both one-hand and two-hand poses using mmWave radar.
+Solutions Radar Sensing View Body Part
+mmMesh [14] AWR1843 Frontal Body
+SynMotion [15] IWR1443 Frontal Body
+mmGPE [16] AWR1843 Frontal Body
+M4esh [17] AWR1843 Frontal Body
+mmBody [18] Arbe Phoenix Frontal Body
+m3Track [19] IWR1443 Frontal Body
+HUPR [20] IWR1843 Frontal Body
+Mi-Mesh [21] AWR1443 Frontal Body
+SUPER [22] IWR6843 Frontal Upper
+mm4Arm [27] IWR6843 Frontal One hand
+mmHand [26] AWR1443 Frontal One hand
+mmHand [25] IWR1843 Frontal One hand
+mmEgo [23] IWR6843 Egocentric Body
+Argus [24] BGT60TR13C $\times$ 2 Egocentric Body
+mmEgoHand (Ours) IWR6843 Egocentric 1 or 2 Hands
 2 Related Work
-2.1 Gesture Recognition with Cameras
-In the field of computer vision, a series of methods have been developed for hand gesture recognition, most of which rely on RGB/D cameras that provide rich geometric information [31, 48].
+2.1 Hand Perception with Cameras
+In the field of computer vision, a series of methods have been developed for hand pose estimation and gesture recognition, most of which rely on RGB/D cameras that provide rich geometric information [32, 33].
 Depth cameras perform excellently in handling complex gestures, such as the effective estimation of gesture joints through the combination of the PointNet model  [34], and Zimmermann’s proposal of the 3D Transformer network to further improve recognition accuracy  [35].
@@
 TMMF  [38] focuses on the single-stage recognition of continuous gestures, emphasizing the importance of temporal information in enhancing the continuity of gesture recognition.
-Figure 2: EgoHand  takes the head-mounted millimeter-wave radar signals and IMU data to generate hand keypoints in the first stage, and further estimates the gesture category in the second stage.
-The camera is used solely for label generation during training and is not involved in inference.
-2.2 Gesture Recognition with mmWave Radars
-Millimeter-wave radar has significant advantages in spatial resolution and penetration ability, providing a reliable and efficient solution for human gesture recognition and pose estimation [26, 24, 30, 32].
+However, these vision-based methods inherently depend on capturing detailed visual information of users’ hands and surrounding environments, which raises privacy concerns in sensitive applications and shared spaces.
+In addition, they may suffer from occlusions in egocentric scenarios, where hands frequently self-occlude or interact with objects (e.g., VR controllers).
+2.2 Hand Perception with mmWave Radars
+In the field of radio frequency (RF) sensing, studies have shown that Wi-Fi and customized radars are capable of estimating the poses of multiple people [3, 4, 29, 39], but the performance in terms of stability and accuracy under complex environments is still lacking, and often requires additional hardware such as increased antenna count and device numbers to enhance performance.
+In contrast, commercial millimeter-wave radar has significant advantages in spatial resolution and penetration ability, providing a reliable and efficient solution for human gesture recognition and pose estimation [9, 7, 40, 41].
 It is noteworthy that recent research has demonstrated the progress of millimeter-wave radar in behavior detection, such as mHomeGes [7] and mTransSee [9], which use millimeter-wave signals to achieve real-time arm gesture recognition and environment-independent gesture recognition, respectively.
 mmASL [41] extracts frequency features from 60GHz millimeter-wave signals and uses a multitask neural network to recognize American Sign Language.
-Pantomime et.al.
+Pantomime et al.
 [40] uses millimeter-wave radar to compute sparse 3D point clouds for gesture recognition in their self-collected dataset.
-However, most of these mmWave radar based systems are designed to be fixed installations; adapting them to a user’s free movement in space, and overcoming limitations of viewpoint and location remain challenges, which may limit their application in VR interactions.
+It can be observed in Table.
+I that most existing methods rely on radars fixed in a static position, constraining users to operate strictly within the radar’s frontal field of view.
+To overcome this limitation, mmEgo [23] and Argus [24] propose mounting radars on the user’s head to enable egocentric body tracking, allowing more natural and unrestricted movements.
+Building on this design, our work extends egocentric millimeter-wave sensing to fine-grained hand pose estimation and hand gesture recognition, broadening the scope of egocentric radar-based interaction.
+Moreover, to the best of our knowledge, our system is the first egocentric mmWave-based solution capable of supporting both one-handed and two-handed interactions, significantly expanding the range of applicable interaction scenarios.
 3 Methods
+The technical details are described below.
 3.1 Data Preprocessing
-In our setup, mmWave radar transmits and receives frequency modulated continuous wave (FMCW) signals chirp by chirp (64 chirps per frame, 20 frames per seconds), and mixes the received signals with the transmitted signals to obtain an intermediate frequency (IF) signal [2, 17, 9].
+In our setup, we use a TI IWR6843 mmWave radar, a 1.75 W low-power radar, to capture hand movements.
+mmWave radar transmits and receives frequency modulated continuous wave (FMCW) signals chirp by chirp (64 chirps per frame, 20 frames per seconds), and mixes the received signals with the transmitted signals to obtain an intermediate frequency (IF) signal [42, 43, 21].
 Due to the sparse point cloud in mmWave radar data, typically consisting of only a few points, hand reflections are inadequately represented, and minor posture changes can cause significant variations, leading to inconsistent representations.
 To address this issue, we adopted a preprocessing method similar to that in mHomeGes [7] to obtain a richer representation of the radar data.
-$\bullet$ Range FFT. We apply range FFT on every IF chirp, which reveals the frequency difference between the received and transmitted signals, to obtain distance heatmaps of reflection objects.
+Figure 2: Millimeter-wave radar signal processing involves several Fourier transform steps focusing on the time domain, chirp signals, and the receiving antenna dimension.
+$\bullet$ Range FFT.
+We apply range FFT on every IF chirp, which reveals the frequency difference between the received and transmitted signals, to obtain distance heatmaps of reflection objects.
+To reduce interference from reflections in the surrounding environment, we retain only the signals within 2 meters after the range FFT, as the distance from the user’s head-mounted mmWave radar to the hands is typically less than 2 meters for most body types.
 $\bullet$ Doppler FFT.
-There exists a phase difference between neighboring chirps caused by the object movements, we use Doppler FFT on chirps in one frame along the direction of the phase change caused by the Doppler effect to estimate the object’s velocity, the outputs called range-Doppler heatmaps.
+There exists a phase difference between neighboring chirps caused by the object movements, we use apply Doppler FFT on chirps in one frame along the direction of the phase change caused by the Doppler effect to estimate the object’s velocity, the outputs called range-Doppler heatmaps.
+The velocity information is illustrated at the horizontal axis and the distance information at the vertical axis in Fig.
+2.
+The Range FFT and Doppler FFT can be formulated as follows:
+$S_{rd}=\mathcal{F}_{chirp}{(\mathcal{F}_{sampling}{(\text{IF})})}$ (1)
+where $S_{rd}$ represents the range-Doppler heatmaps; $\mathcal{F}_{sampling},\mathcal{F}_{chirp}$ signifies the FFT pertaining to the sampling point dimension and chirp dimension, respectively.
+Figure 3: mmEgoHand  takes the head-mounted millimeter-wave radar signals and IMU data to generate hand keypoints.
+The camera is used solely for label generation during training and is not involved in inference.
 $\bullet$ Angle FFT.
@@
 We use angle FFT is executed on the range-Doppler heatmaps along the receiving antenna dimension to acquire range-angle heatmaps, which responds to the distance and angle information of the reflection objects.
-We concatenate range-Doppler heatmaps and range-angle heatmaps along the vertical axis for the deep network input, as shown in Fig. 3. In contrast, we do not preprocess IMU series.
-Figure 3: Millimeter-wave radar signal processing involves several Fourier transform steps focusing on the time domain, chirp signals, and the receiving antenna dimension.
+The angle information is illustrated at the horizontal axis and the distance information at the vertical axis in Fig.
+2.
+The process can be formulated as follows:
+$S_{ra}=\mathcal{F}_{rx}{(S_{rd})}$ (2)
+where $S_{ra}$ denotes the range-angle heatmaps, and $\mathcal{F}_{rx}$ refer to the FFT operation along the receiving antenna dimension.
+We concatenate range-Doppler heatmaps and range-angle heatmaps along the vertical axis for the deep network input, as shown in Fig. 2.
 3.2 Deep Network Design
-We use an 15.85 mW low-power mmWave radar to capture hand movements.
+We use TI IWR6843 mmWave radar to capture hand movements.
 However, the radar signal also contains noise from head movements.
@@
 Therefore, we use a Transformer-based structure for representation learning.
-Another design choice arose from experimental results: directly predicting gesture classes led to poor performance, with accuracy around 65%.
-We therefore opted to first generate an intermediate representation as a transition step, then use it to classify gestures.
-Research using cameras [8, 16] has shown that hand skeletons serve as a promising and interpretable intermediate representation, so we adopted this approach.
-The third design consideration is handling gestures involving different numbers of hands: 0, 1, or 2 hands may be present during interactions.
-Unlike OCHID-Fi [46], which assumes a fixed output of keypoints for one hand, we require an approach that can adapt to varying numbers of hands.
-This is similar to the object detection task, where the model must flexibly output coordinates for a variable number of targets.
-Therefore, we adapt DETR [6], an end-to-end object detection framework, to regress the coordinates of hand keypoints.
-Based on the above considerations, we propose the framework shown in Fig. 2 and the network structure illustrated in Fig. 4. More detailed operations will be introduced next.
-Figure 4: mmVR deep network consists of four main novel components. (a) mmWave Radar Encoder, (b) IMU Encoder, (c) Pose Decoder, and (d) Context Decoder.
+Another consideration is handling poses involving different numbers of hands: 0, 1, or 2 hands may be present during interactions.
+Unlike mmHand [25, 26], which assumes a fixed output of keypoints for one hand, our system must dynamically adapt to variable hand counts.
+This is similar to the multi-person pose estimation task, where the model must flexibly output coordinates for a variable number of people.
+To address this, we adapt the Person-in-WiFi 3D framework [29], an end-to-end solution for multi-person pose estimation, to regress hand keypoints for multiple hands.
+Guided by these insights, we propose the framework shown in Fig.
+3 and the network structure illustrated in Fig.
+4.
+$\bullet$ Input and Output.
+We take a 2-second sequence of mmWave radar and IMU signals as input and output the corresponding 3D hand pose sequence.
+The mmWave input is represented as $x_{mm}\in\mathbb{R}^{30\times 256\times 128}$, and the IMU input as $x_{imu}\in\mathbb{R}^{30\times 2\times 3}$, where 30 is the number of sampled frames.
+The output is a 3D keypoint sequence of shape $\mathbb{R}^{30\times h\times K\times 3}$, where $h$ is the number of hands, and $K$ is the number of keypoints per hand.
+To align the modalities temporally, we divide the 2-second window into 30 uniform patches.
+Each mmWave patch is flattened into a vector of shape $\mathbb{R}^{32,768\times 1}$ (i.e., $256\times 128$), and each IMU patch—consisting of one sample across 6 channels (3-axis acceleration and 3-axis angular velocity)—is flattened into a vector of shape $\mathbb{R}^{6\times 1}$.
+These patch vectors are then independently fed into an mmWave encoder and an IMU encoder to extract modality-specific temporal features for subsequent fusion.
+Figure 4: mmEgoHand  deep network consists of four main novel components. (a) mmWave Radar Encoder, (b) IMU Encoder, (c) Pose Decoder, and (d) Context Decoder.
 $\bullet$ mmWave Radar Encoder.
 This encoder integrates six encoder blocks to process mmWave radar streams for mmWave radar embeddings.
-Each block comprises a multi-head self-attention module and a Feed-Forward Network (FFN), which are fundamental components of the Transformer architecture  [35].
+Each block comprises a multi-head self-attention module and a Feed-Forward Network (FFN), which are fundamental components of the Transformer architecture [46], shown in Fig.
+4 (a).
 $\bullet$ IMU Encoder.
-This part of the architecture is similar to the mmWave Radar Encoder, which is used for IMU data embeddings.
-The encoded IMU embeddings are concatenated with the mmWave radar embeddings for further Pose Decoder and Context Decoder.
+This encoder is similar to the mmWave Radar Encoder, which is used for IMU data embeddings.
+The encoded IMU embeddings are concatenated with the mmWave radar embeddings for further Pose Decoder and Context Decoder, shown in Fig.
+4 (b).
 $\bullet$ Pose Decoder.
-The structure of the decoder layer is the basic module of the DETR [6] decoder, and we stacked three such layers.
-One Pose Decoder block contains one multi-head self-attention module and one multi-head cross-attention module.
+The structure of the decoder layer is the basic module of the Person-in-WiFi 3D [29] decoder, and we stacked three such layers.
+As shown in Fig.
+4 (c), one Pose Decoder block contains one multi-head self-attention module and one multi-head cross-attention module.
 The multi-head cross-attention module compute attention matrix between Pose Queries ($Q_{pose}=[q_{1},q_{2},\ldots,q_{C}]\in\mathbb{R}^{C\times D}$) and the fused embeddings ($E_{fused}$) from mmWave Radar Encoder and IMU Encoder.
@@
 $X_{pose}=\text{PoseDecoder}(Q_{pose},E_{fused})$ (3)
+where $E_{fused}$ represents the fused features from mmWave radar embeddings and IMU embeddings.
 $\bullet$ Context Decoder.
-One Context Decoder block includes one multi-head self-attention module and two multi-head cross-attention modules, and we stack 30 such blocks in the Context Decoder.
+As shown in Fig.
+4 (d), one Context Decoder block includes one multi-head self-attention module and two multi-head cross-attention modules, and we stack 30 such blocks in the Context Decoder.
 Context Decoder learns the globally temporal relationships among different frames to refine the hand keypoint estimation.
-Specifically, Context Decoder utilizes the multi-head cross-attention to compute cross-attention matrix between the current query and previous pose features from Pose Decoder, iteratively.
+Specifically, Context Decoder utilizes the multi-head cross-attention to compute a cross-attention matrix between the current query and previous pose features from Pose Decoder, iteratively.
 At last, two Multi-Layer Perceptrons (MLPs) are respectively to generate the keypoints of the hand candidates and the corresponding confidence scores indicating the authenticity of these hand candidates.
@@
 In our settings, $C=100$, $K=21$, and $n=30$.
+The Context Decoder consists of 30 Transformer blocks, each responsible for predicting the pose at a single time step.
+Each block takes as input the current frame along with the refined representation from the previous frame, enabling cross-frame temporal modeling through cross-attention.
+This dual-decoder design explicitly decomposes the egocentric hand tracking challenge: the Pose Decoder specializes in spatial joint localization by resolving per-frame ambiguities through cross-attention between pose queries $Q_{pose}$ and sensor embeddings $E_{fused}$; conversely, the Context Decoder captures temporal gesture kinematics via iterative feature refinement across frames, enabling robust recognition of dynamic motions.
+This separation aligns with the observation that spatial precision and temporal coherence require specialized processing under head-motion interference.
 $\bullet$ Loss Function. We implement a Hungarian Matching algorithm [30] to ensure a unique prediction for the ground-truth 3D hand keypoints of each hand candidate.
 $L_{kpt}=\text{HungarianMatch}(H_{set},H_{gt})$ (5)
-where $H_{gt}$ denotes the ground-truth hand keypoint coordinates.
-$H_{set}$ is selected up to two hand candidates based on $H_{score}$.
-If there are two hands, their keypoints are concatenated and fed into the recognition network.
-If there is only one hand, it is duplicated before concatenation, ensuring that the recognition network has a fixed input size.
-Further, we apply a cross-entropy function to compute the gesture recognition loss.
-$L_{cls}=\text{CrossEntropy}(y,\hat{y})$ (4)
-where $y$ and $\hat{y}$ represent the gesture label and recognition confidence score, respectively.
-The total loss to optimize the deep network is a weighted combination of hand keypoint regression loss and the gesture regression loss.
-$Loss=L_{kpt}+\lambda L_{cls}$ (5)
-where $\lambda$ is a weight parameter balancing the hand keypoint regression and gesture recognition.
-$\bullet$ Training Details.
-We employed the Adam optimizer [19] to train deep network on a NVIDIA 3090 GPU.
+where $H_{gt}$ denotes the ground-truth hand keypoint coordinates. $H_{set}$ is selected up to two hand candidates based on $H_{score}$.
+$\bullet$ Implementation Details.
+We employed the Adam optimizer [47] to train a deep network on an NVIDIA 3090 GPU.
 The batch size was 32.
-The training was carried out 500 epochs.
+The training was carried out for 500 epochs.
 The initial learning rate was 0.001 and was halved every 100 epochs.
 At the start of each epoch, the training samples were randomly shuffled.
-To balance the keypoint regression and gesture recognition, we experimentally set $\lambda=1$ in Equation.
+The resulting model contains 55.42M parameters, with an inference GPU memory of 657 MB and an end-to-end inference time of 42.03ms for 30 frames on the RTX 3090 GPU.
+These results demonstrate that our framework already achieves real-time performance.
+$\bullet$ Downstream gesture recognition task.
+The output of mmEgoHand  can be seamlessly applied to downstream tasks such as gesture recognition.
+Given a sequence of predicted hand poses $H_{set}$, we feed it into a gesture classification network (e.g., a GNN or ResNet) to identify the performed gesture.
+To ensure a consistent input size for the recognition model, we handle single- and dual-hand scenarios differently: if two hands are present, their keypoints are concatenated directly; if only one hand is detected, its keypoints are duplicated before concatenation.
+This design allows the recognition network to operate on a fixed-size input regardless of the number of hands.
+4 Experiments
+4.1 Data Acquisition
+Data was collected under the approval of the IRB.
+Figure 5: Hardware setup.
+The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, including three postures in three scenes, i.e., standing at scene #3, lying at scene #2, and sitting at scene #1, respectively.
+(1) Hardware Configurations.
+The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, shown in Fig.
 5.
-4 Experiments
-4.1 Dataset Acquisition
-Dataset was collected under the approval of IRB.
-(1) Hardware Configurations. The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, shown in Fig. 5. These devices were manually synchronized before recording started.
+The millimeter-wave radar and IMU device are mounted on a plastic bracket and secured to the head using an adjustable strap.
+A tripod is used to position the camera for hand motion capture and hand pose labeling.
+All three devices are controlled and synchronized via a desktop computer, ensuring precise temporal synchronization during data recording.
 $\bullet$ Millimeter-wave radar.
 One TI IWR6843ISK radar generates frequency modulated continuous wave (FMCW) signals at frequencies of 60GHz-64GHz.
-Its transmit power is only 15.85 mW, slightly lower than typical Wi-Fi devices, making it safe for human exposure.
+Its power consumption is only 1.75W, slightly lower than typical Wi-Fi devices, making it safe for human exposure.
 The radar has three transmitting antennas and four receiving antennas.
 The transmit parameters are set to 20 frames per second with 64 chirps per frame.
-The analog-to-digital converter (ADC) sampling rate was 256, and the raw radar data was obtained by mounting the radar on a TI DCA1000EVM board.
+The analog-to-digital converter (ADC) sampling rate is 256, and the raw radar data is obtained by mounting the radar on a TI DCA1000EVM board.
 The raw radar data is $\in 20t\times 3\times 4\times 256\times 64$, where $t$ is the time in seconds.
-The range-Doppler and range-angle heatmaps are obtained after the preprocessing process, both of which are $\in 20t\times 256\times 64$.
+The range-Doppler and range-angle heatmaps are obtained after the preprocessing process described in Sec.
+3.1, both of which are $\in 20t\times 256\times 64$.
+These heatmaps are then combined along the distance dimension to obtain model inputs $\in 20t\times 256\times 128$.
 $\bullet$ IMU.
@@
 (2) Dataset diversity.
-We recruited 10 volunteers with heights ranging from 161 to 182 centimeters and weights between 44 to 73 kilograms.
-We conducted our experiment across three different scenes and three distinct postures, i.e., sitting, lying, and standing, wherein each volunteer engaged in 8 types of human-computer interaction gestures.
-These gestures include click, swipe leftward, swipe rightward, swipe upward, swipe downward, swipe to bottom right, zoom in, and zoom out.
-The latter two are two-handed gestures and the rest are single-handed gestures.
-The selection of these gestures and postures is inspired by the Apple Vision Pro, such as confirming selection and scrolling up the page, which has demonstrated that this set of gestures can facilitate a wide range of VR interactive functions, as shown in the promotional videos released by Apple [^1] Youtube link and Bilibili link.
-In each scene and posture, volunteers were instructed to perform each gesture 20 times, with each gesture completed within 2 seconds.
-Of the 10 volunteers, two are left-handed, and in performing single-handed gestures, they have been assigned tasks to execute using their left and right hands, respectively.
-The remaining eight volunteers are right-handed, among whom six have been arranged to carry out interactive gestures using their dominant right hand, while two have been assigned to perform gestures using their non-dominant left hand.
-Thus, there are $2\times 2+6+2=12$ different settings for the left and right hand executing gestures in total.
+We recruited 10 volunteers, whose heights ranged from 161 to 182 centimeters and weights from 44 to 73 kilograms.
+The experiment was conducted across three different environments and under three distinct postures: sitting, lying, and standing.
+In each condition, the participants performed eight types of human-computer interaction gestures: click, swipe left, swipe right, swipe up, swipe down, swipe to the bottom-right, zoom in, and zoom out.
+Among these, the last two are two-handed gestures, while the rest are performed with a single hand.
+The selection of gestures and postures was inspired by the Apple Vision Pro, which employs similar interactions—such as confirming selections or scrolling through content—for a broad range of VR functionalities.
+In every scene and posture, each volunteer was asked to perform each gesture 20 times, with each gesture completed within a 2-second duration.
+Among the 10 volunteers, two are left-handed, and in performing single-handed gestures, they were instructed to perform gestures using their left and right hands, respectively.
+The remaining eight volunteers are right-handed; six of them performed gestures with their dominant right hand, while the other two were assigned to use their non-dominant left hand.
+Consequently, there are $2\times 2+6+2=12$ different settings for the left and right hand executing gestures in total.
 Table.
-1 shows this experimental setting.
+II summarizes this experimental configuration.
 This setup has been designed for scenarios where a user’s preferred hand is occupied, ensuring that the system can still recognize and respond to gestures executed by the non-preferred hand.
+TABLE II: We recruited 10 volunteers to perform VR interaction gestures.
+They were asked to execute gestures by their preferred hand or non-preferred hand.
+Pre.Hand for Preferred hand; Exe.Hand for Executing Hand; H.
+for Height; W.
+for Weight.
 No. ID Pre.Hand Exe.Hand Gender H.(cm) W.(kg) BMI
@@
 12 10 Right Left Female 163 48 18.07
-Table 1: We recruited 10 volunteers to perform VR interaction gestures.
-They were asked to execute gestures by their preferred hand or non-preferred hand.
-Pre.Hand for Preferred hand; Exe.Hand for Executing Hand; H.
-for Height; W.
-for Weight.
-Figure 5: Hardware setup.
-The data collection hardware consists of a millimeter-wave radar, an IMU, and a camera, including three postures in three scenes, i.e., standing at scene #3, lying at scene #2, and sitting at scene #1, respectively.
 (3) Dataset statistics.
-The dataset contains 5760 (3 scenes $\times$ 12 hand execuring settings $\times$ 8 gestures $\times$ 20 repetitions) gesture samples, including synchronized data of millimeter-wave radar, IMUs, and videos.
-We uniformly select the 5th, 10th, 15th, and 20th repetitions for test, and the remaining 16 repetitions for training.
-We utilize Google MediaPipe Hand Landmark SDK to extract 3D hand keypoints from videos as the ground truth for deep network training and evaluation.
-We observe that the SDK sometimes misses to extract hand keypoints, especially in situations when hands move very fast.
-If the SDK misses 40% or more frames in a particular video, we discard it.
-Otherwise, we downsample or upsample the hand keypoints to 30 frames per gesture sample.
-Although MediaPipe’s 3D positions may lack high precision, this approach significantly simplifies the annotation process.
-Consequently, we have a dataset that consists of 5206 gesture samples, detailed statistics shown in Table.
-2.
+The dataset comprises a total of 5,760 gesture samples, collected across 3 scenes, 12 hand-executing settings, 8 gesture types, and 20 repetitions per gesture (3×12×8×20).
+Each sample includes synchronized data from millimeter-wave radar, IMUs, and video recordings.
+For data partitioning, we uniformly select the 5th, 10th, 15th, and 20th repetitions for testing, and use the remaining 16 repetitions for training.
+To provide supervision for deep network training and evaluation, we employ the Google MediaPipe Hand Landmark SDK to extract 3D hand keypoints from the video data.
+However, we observe that the SDK occasionally fails to detect keypoints, particularly when the hand moves rapidly.
+If more than 40% of frames in a video are missing keypoints, the corresponding sample is discarded.
+Otherwise, the extracted keypoints are temporally downsampled or upsampled to 30 frames per gesture sample to maintain consistency.
+Although MediaPipe’s 3D keypoints may not offer high-precision tracking, this automated approach significantly reduces the annotation burden.
+After filtering out invalid samples, the final dataset contains 5,206 gesture instances.
+Detailed statistics are presented in Table III.
+TABLE III: Dataset samples. Volunteers executed the gestures with postures of sitting in scene #1, lying in in scene #2, and standing at scene #3, respectively.
 sitting in s#1 lying in s#2 standing in s#3 all
@@
 #test 350 351 353 1054
-Table 2: Dataset samples. Volunteers executed the gestures with postures of sitting in scene #1, lying in in scene #2, and standing at scene #3, respectively.
 4.2 Evaluation Metrics
@@
 (2) We also adopt gesture recognition Accuracy, Precision, Recall, and F1-score for the evaluation.
-4.3 In-domain Results
-Overall performance.
-Table.
-3 presents the main results of EgoHand  under the data splitting strategy shown in Table.
-2.
-The one-stage approaches directly output the gesture classification results without regressing the hand keypoints.
-In contrast, EgoHand  outputs hand keypoints as the intermediate features and conduct gesture recognition over the keypoints.
-We evaluate current mainstream and basic classification networks.
-Table.
-3 demonstrates that EgoHand  outperforms one-stage approaches by a large margin (except EgoHand  (ViT [10])).
-This result demonstrates that the novel end-to-end Transformer architecture proposed in our first stage is highly effective.
-A more intriguing finding is that, despite not achieving particularly good MPJPE in the first stage, intermediate hand keypoint positions significantly boost the accuracy of gesture recognition.
-This suggests that future work could consider promoting hand keypoint estimation to enhance gesture recognition.
-In light of the good performance of ResNet50 in classification tasks, this network has been selected as the gesture recognition backbone.
-Metrics MPJPE Acc(%) Precision Recall F1
+TABLE IV: mmEgoHand outperforms one-stage methods such as mGesNet [7], mSeeNet [9], and mmGesture [31], by a large margin.
+MPJPE is measured in millimeters(mm).
+mm4Arm [27] is a single-hand pose estimation method, trained and evaluated solely on single-hand data.
+Compared to simultaneous detection of both single-hand and two-hand poses, this is a considerably simpler task.
+Nevertheless, our method significantly outperforms mm4Arm despite addressing the more challenging setting.
+[…]
+mmEgoHand (mmWave+IMU) 72.73mm / / / /
+mmEgoHand (mmWave) 96.42mm / / / /
+mmEgoHand (mmWave+IMU, w/o Context Decoder) 109.02mm / / / /
+mm4Arm [27] (mmWave+IMU) 165.19mm / / / /
+two-stage (mmEgoHand+ResNet50 [48]) / 90.80 93.34 93.20 93.19
+two-stage (mmEgoHand+LSTM) / 89.26 91.56 91.48 91.46
+two-stage (mmEgoHand+GCN) / 85.36 87.16 87.12 87.09
+two-stage (mmEgoHand+ViT [49]) / 77.95 81.39 80.19 79.28
+one-stage (mSeeNet [9]) / 84.60 86.73 86.59 86.60
+one-stage (mGesNet [7]) / 81.34 81.98 81.56 81.64
+one-stage (mmGesture [31]) / 68.66 70.01 68.83 68.81
 one-stage (LSTM) / 67.99 70.58 68.35 67.80
-one-stage (ViT) / 59.66 67.12 59.77 60.75
-one-stage (ResNet50) / 60.09 61.79 61.19 60.64
-one-stage (mGesNet) / 81.34 81.98 81.56 81.64
-one-stage (mSeeNet) / 84.60 86.73 86.59 86.60
-one-stage (mmGesture) / 68.66 70.01 68.83 68.81
-EgoHand (GCN) 81.95 85.36 87.16 87.12 87.09
-EgoHand (LSTM) 75.70 89.26 91.56 91.48 91.46
-EgoHand (ViT) 91.20 77.95 81.39 80.19 79.28
-EgoHand (ResNet50) 72.73 90.80 93.34 93.20 93.19
-Table 3: EgoHand outperforms one-stage methods such as mGesNet [24], mSeeNet [26], and mmGesture [40], by a large margin. MPJPE is measured in millimeters(mm).
+one-stage (ResNet50 [48]) / 60.09 61.79 61.19 60.64
+one-stage (ViT [49]) / 59.66 67.12 59.77 60.75
+4.3 Results
+(1) Overall performance.
+$\bullet$ IMU works.
+Table IV presents a comprehensive evaluation of mmEgoHand’s performance in both hand pose estimation and downstream gesture recognition, following the data partitioning strategy detailed in Table III.
+Our system achieves a Mean Per Joint Position Error (MPJPE) of 72.73mm, demonstrating a 24.6% improvement over mmWave-only approaches (96.42mm MPJPE).
+This significant enhancement not only confirms the presence of head motion in our data collection process but also validates the effectiveness of our IMU signal fusion strategy for egocentric hand pose estimation.
+$\bullet$ Context Decoder works.
+To further assess the contribution of the Context Decoder, we conducted a controlled experiment by training the model with only the Pose Decoder.
+In this configuration, the system achieved an MPJPE of 109.92mm.
+These results confirm that the Context Decoder significantly enhances performance by modeling sequential context across frames.
+$\bullet$ SOTA Pose Estimation.
+Besides, due to the current limitations of mmHand and mm4Arm—namely, their support for single-hand estimation only and their closed-source implementations [25, 26, 27]—we independently reimplemented mm4Arm [27] and trained it exclusively on single-hand data for hand pose estimation, achieving an MPJPE of 165.19mm.
+This single-hand setting is inherently simpler than our scenario, which simultaneously estimates both single- and two-hand poses.
+Nevertheless, our method significantly outperforms mm4Arm despite addressing this more challenging task.
+$\bullet$ SOTA Gesture Recognition.
+Furthermore, we use the hand poses estimated by mmEgoHand’s intermediate features and feed them into simple classification models such as ResNet [48], LSTM, GCN, and ViT [49] for gesture recognition, as highlighted in blue in TableIV.
+All models achieve strong performance under this two-stage pipeline.
+For instance, using ResNet-50 yields a gesture recognition accuracy of 90.80% and an F1-score of 93.19.
+These results are 20–30 percentage points higher than directly applying LSTM, ResNet-50, or ViT to raw mmWave and IMU data.
+Moreover, mmEgoHand  also outperforms carefully designed state-of-the-art methods for mmWave gesture recognition, such as mGesNet [7], mSeeNet [9], and mmGesture [31].
+This finding suggests that future research could benefit from incorporating hand pose estimation as an intermediate representation to enhance gesture recognition performance.
+TABLE V: Accuracy, Precision, Recall and F1 for each gesture.
 Gesture Acc(%) Precision Recall F1
@@
 mean 90.80 93.34 93.20 93.19
-Table 4: Accuracy, Precision, Recall and F1 for each gesture.
-Figure 6: EgoHand  hand pose estimation examples, showing click, swipe downward, zoom in, and swipe leftward, respectively. (Images here are only used for visualization).
-PJPEs for fingers.
-In addition, PJPEs of the thumb finger, index finger, middle finger, ring finger, and little finger are 60.18, 70.10, 76.42, 84.44, and 91.79, respectively.
-The thumb and index finger exhibit the least prediction errors, whereas the largest errors are found in the ring and little fingers.
-This is because, from the perspective of a mmWave radar worn on the head, the thumb and index finger are the most discernible, while the other fingers tend to be frequently obscured by the thumb and index finger.
-Visualization.
+(2) PJPEs for fingers.
+In addition, the PJPEs of the thumb, index, middle, ring, and little fingers are 60.18, 70.10, 76.42, 84.44, and 91.79, respectively.
+The thumb and index fingers exhibit the smallest prediction errors, while the ring and little fingers show the largest.
+This is because, from the perspective of a head-mounted mmWave radar, the thumb and index fingers are more prominently exposed, whereas the other fingers are often occluded by them, making accurate estimation more challenging.
+Nevertheless, thanks to the penetration capability of mmWave signals [23], our model is still able to achieve reasonable estimation performance for these occluded fingers.
+Figure 6: mmEgoHand  hand pose estimation examples, showing click, swipe downward, zoom in, and swipe leftward, respectively. (Images here are only used for visualization).
+(3) Visualization.
 Fig.
-6 illustrates the visualization of hand keypoint predictions by the EgoHand  for various gestures, showcasing different volunteers performing gestures such as click, swipe downward, zoom in, and swipe leftward.
-Each set of gestures is depicted in three rows: the first row presents the video frames captured during the gestures, the second row shows the hand keypoints derived from the video frames using the Google SDK, and the third row displays the predicted hand keypoints output by EgoHand.
-It can be seen that EgoHand  can automatically detect the number of hands, whether they are left or right hands, as well as the movement of hand joints throughout the gesture executing process, which effectively alleviates the difficulty of subsequent gesture recognition.
-Gesture recognition.
-Table.
-4 shows the Accuracy, Precision, Recall and F1 Score for each gesture.
-We can see that the accuracy of single-hand gestures is generally above 90%, whereas the accuracy of two-hand gestures is slightly lower but still remains above 80%.
-The most dominant classification errors are the mutual misclassifications between the two-hand gestures of zoom in and zoom out.
-This is because, during the data collection, volunteers have no strict guidance on which diagonal direction should be used for zoom in and zoom out, introducing more variability compared to single-hand gestures.
-Besides, The MPJPEs of zoom in and zoom out in the first stage are generally higher than the single-hand gestures, leading to suboptimal performance in the subsequent gesture classification.
-Modality ablation study.
-Table.
-5 shows the experiments relating to the impact of IMU data in EgoHand.
-The experiments employed strategies using only millimeter-wave radar data, only IMU data, and both data streams.
-Results indicate that using only IMU data to estimate gestures results in weakest performance in terms of both mean MPJPE and gesture recognition accuracy.
-Relying only on mmWave radar data also leads to a decrease in metrics compared to the both using millimeter-wave radar and IMU data.
-This experiment demonstrates that data captured by IMUs can effectively compensate for human body or head movements, thereby enhancing the accuracy of hand keypoint estimation as well as gesture recognition.
-4.4 Cross-domain Evaluation
-Cross-person.
-We utilize two cross-person validation approaches to evaluate the performance of EgoHand  in gesture recognition on unseen users.
-(1) We use dataset from No.1 to No.8 listed in Table.
-1 as the training set to train EgoHand, and test the trained EgoHand  with dataset from No.9 to No.12.
-(2) We select one of 12 rows from Table.
-1 for testing and use the remaining rows for training; and we iteratively repeat this process 12 times and ultimately compute the average value.
-Table.
-6 shows that EgoHand  maintains a certain level of recognition capability for unseen persons, achieving a gesture recognition accuracy of 63.11 and 61.96% in the above two training/test splitting strategy, respectively.
-Moreover, the gesture recognition accuracy significantly improves by 10%-20% with one-shot or two-shot fine-tuning, using only one or two samples from the target person/s to fine-tune EgoHand.
-Input MPJPE Acc(%) Precision Recall F1
-IMUs 120.18 68.80 70.13 69.80 69.36
-mmWave 96.42 86.04 88.70 88.55 88.54
-Fusion 72.73 90.80 93.34 93.20 93.19
-Table 5: With mmWave radar and IMUs, EgoHand  performs the best. MPJPE is measured in millimeters(mm).
+6 illustrates the visualization of hand keypoint predictions by the mmEgoHand  for various gestures, showcasing different volunteers performing gestures such as click, swipe downward, zoom in, and swipe leftward.
+Each set of gestures is depicted in three rows: the first row presents the video frames captured during the gestures, the second row shows the hand keypoints derived from the video frames using the Google SDK, and the third row displays the predicted hand keypoints output by mmEgoHand.
+It can be seen that mmEgoHand  can automatically detect the number of hands, whether they are left or right hands, as well as the movement of hand joints throughout the gesture executing process, which effectively alleviates the difficulty of subsequent gesture recognition.
+Figure 7: The confusion matrix of gesture recognition.
+(4) Downstream gesture recognition.
+Fig.
+7 presents the confusion matrix for recognizing eight human-computer interaction gestures.
+Each cell in the matrix indicates both the percentage and the absolute number of samples classified into each category.
+Table V further reports the Accuracy, Precision, Recall, and F1 Score for each gesture.
+We observe that single-hand gestures generally achieve high accuracy, often exceeding 90%, while two-hand gestures exhibit slightly lower performance, though still above 80%.
+The most prominent recognition errors occur between the two-hand gestures zoom in and zoom out, which are frequently misclassified with each other.
+This confusion is largely attributed to the lack of strict instructions given to participants during data collection regarding the specific diagonal direction of movement for these gestures, resulting in greater intra-class variability.
+Additionally, we observe that the MPJPE for zoom in and zoom out is notably higher compared to single-hand gestures, contributing to reduced accuracy in the downstream gesture classification task.
+4.4 Few-shot Gesture Recognition
+We further evaluate the generalization ability of mmEgoHand for gesture recognition under few-shot fine-tuning in challenging real-world scenarios, including cross-person, cross-posture, and cross-hand settings.
+Please note that the primary goal of this evaluation is to demonstrate the baseline generalization capability of our framework.
+Achieving stronger generalization performance would require incorporating additional techniques such as domain adaptation and domain generalization.
+To facilitate future research in this direction, we release our dataset publicly to support further exploration of these advanced generalization methods.
+(1) Cross-person.
+We adopt two cross-person validation strategies to evaluate the gesture recognition performance of mmEgoHand  on unseen users.
+(1) In the first strategy, we use data from No.1 to No.8 (as listed in Table.
+II) for training, and test the model on data from No.9 to No.12.
+(2) In the second strategy, we perform leave-one-subject-out evaluation: in each iteration, data from one row in Table.
+II is used for testing, while data from the remaining rows is used for training.
+This process is repeated for all 12 rows, and the average performance is reported.
+As shown in Table VI, mmEgoHand demonstrates a certain degree of generalization ability to unseen users, achieving gesture recognition accuracies of 63.11% and 61.96% under the two strategies, respectively.
+Furthermore, the accuracy improves significantly—by 10% to 20%—when one-shot or two-shot fine-tuning is applied, using only one or two labeled samples from the target user to adapt the mmEgoHand.
+TABLE VI: Gesture recognition tests on unseen volunteers.
+mmEgoHand  maintains a certain level of recognition capability for unseern persons, and can be improved by 10%-20% with one-shot or two-shot fine-tuning.
 #Training; #Test #-Shot Acc(%) Precision Recall F1
 zero 63.11 64.42 63.17 62.74
-No.1-8; No.9-12 (Table. 1) one 73.99 74.50 73.60 73.57
+No.1-8; No.9-12 (Table. II) one 73.99 74.50 73.60 73.57
 two 83.97 84.13 83.88 83.87
@@
 two 74.65 77.86 76.51 75.09
-Table 6: Gesture recognition tests on unseen volunteers.
-EgoHand  maintains a certain level of recognition capability for unseern persons, and can be improved by 10%-20% with one-shot or two-shot fine-tuning.
-Figure 7: Gesture recognition accuracy under leave-one-person-out setting.
+Figure 8: Gesture recognition accuracy in leave-one-person-out setting.
+Fig.
+8 demonstrates the leave-one-person-out results.
+It also shows that mmEgoHand  maintains a certain level of recognition capability and can be improved further by few-shot fine-tuning.
+There is one exception observed with No.
+7, which exhibits a significantly lower accuracy rate.
+Upon reviewing the video footage, we find that this particular volunteer performs the gestures with smaller and more casual movements, deviating considerably from the patterns exhibited by others.
+Consequently, during actual system deployment, what we need is to implement a pre-use calibration protocol, requiring users to calibrate each gesture twice or more before commencing usage.
+(2) Cross-posture and Cross-scene.
+We use data from two postures (or scene) in Table.
+III to train mmEgoHand and tests the trained mmEgoHand on data of the last postures (or scene).
+Similar to the cross-person performance, mmEgoHand  maintains a certain level of recognition capability for unseen postures/scenes, achieving a gesture recognition accuracy of 54.35%, 51.42%, and 51.41% in three cross-posture evaluation, respectively.
+Moreover, the gesture recognition accuracy significantly improves by 20% with one-shot or two-shot fine-tuning.
+This experiment further substantiates the efficiency of pre-use calibration strategy discussed in the above cross-person evaluation.
+TABLE VII: Gesture recognition tests on unseen posture/scene.
+mmEgoHand  maintains a certain level of recognition capability for unseen posture/scene, and can be improved by 20% with one-shot or two-shot fine-tuning.
 #Training; #Test #-Shot Acc(%)
-sitting in scene #1 (in-domain result dataset split) / 87.60
-lying in scene #2 (in-domain result dataset split) / 87.96
-standing in scene #3 (in-domain result dataset split) / 95.92
+sitting in scene #1 (dataset split in Table III) / 87.60
+lying in scene #2 (dataset split in Table III) / 87.96
+standing in scene #3 (dataset split in Table III) / 95.92
 zero 54.35
@@
 two 74.51
-Table 7: Gesture recognition tests on unseen posture/scene.
-EgoHand  maintains a certain level of recognition capability for unseen posture/scene, and can be improved by 20% with one-shot or two-shot fine-tuning.
-Fig.
-7 demonstrates the leave-one-person-out results.
-It also shows that EgoHand  maintains a certain level of recognition capability and can be improved further by few-shot fine-tuning.
-There is one exception observed with No.
-7, which exhibits a significantly lower accuracy rate.
-Upon reviewing the video footage, we find that this particular volunteer performs the gestures with smaller and more casual movements, deviating considerably from the patterns exhibited by others.
-Consequently, during actual system deployment, what we need is to implement a pre-use calibration protocol, requiring users to calibrate each gesture twice or more before commencing usage.
-Cross-posture and Cross-scene.
-We use data from two postures (or scene) in Table.
-2 to train EgoHand and tests the trained EgoHand on data of the last postures (or scene).
-Similar to the cross-person performance, EgoHand  maintains a certain level of recognition capability for unseen postures/scenes, achieving a gesture recognition accuracy of 54.35%, 51.42%, and 51.41% in three cross-posture evaluation, respectively.
-Moreover, the gesture recognition accuracy significantly improves by 20% with one-shot or two-shot fine-tuning.
-This experiment further substantiates the efficiency of pre-use calibration strategy discussed in the above cross-person evaluation.
+TABLE VIII: Gesture recognition tests on unseen hands.
+mmEgoHand  maintains a weak level of recognition capability for unseern hands, and can be largely improved by 30%-50% with one-shot or two-shot fine-tuning.
 Data splitting strategy #Training Hand #Test Hand #-Shot Acc(%)
-In-domain results accuracy of left hand / 90.96
+Table III accuracy of left hand / 90.96
 accuracy of right hand / 93.04
@@
 two 77.18
-Table 8: Gesture recognition tests on unseen hands.
-EgoHand  maintains a weak level of recognition capability for unseern hands, and can be largely improved by 30%-50% with one-shot or two-shot fine-tuning.
-Cross-hand.
+(3) Cross-hand.
 Table.
-8 shows that EgoHand  achieves that the accuracy for each hand exceeds 90% if following the data splitting strategy in in-domain results section.
-We also conduct cross-hand evaluation, training EgoHand  with data of right/left executing hand in Table.
-1, and test the trained EgoHand  with data of left/right executing hand, respectively.
-It is noteworthy that this training strategy also encompasses cross-person challenges.
+VIII shows that mmEgoHand  achieves that the accuracy for each hand exceeds 90% if following the data splitting strategy (Table III).
+We also conduct cross-hand evaluation, training mmEgoHand  with data of right/left executing hand in Table.
+II, and test the trained mmEgoHand  with data of left/right executing hand, respectively.
+Note that this training strategy also encompasses cross-person challenges.
 For example, as shown in Table.
-1, instances from No.6 to No.11 introduce cross-person challenges when the left-hand execution data is used for training and the right-hand execution data is employed for testing.
-Thus, this is a highly challenging data splitting scheme, with Table.
-8 indicating that EgoHand  struggles greatly under such split.
+II, instances from No.6 to No.11 introduce cross-person challenges when the left-hand execution data is used for training and the right-hand execution data is employed for testing.
+Thus, this is a highly challenging data splitting scheme.
 However, upon fine-tuning with just one to two samples, gesture recognition accuracy sees a remarkable improvement of 30-50%, up to 73.53% and 77.18%.
-This indicates that during the training of EgoHand, we should simultaneously provide samples from both the left and right hands.
+This indicates that during the training of mmEgoHand, we should simultaneously provide samples from both the left and right hands.
 5 Conclusion
-In this paper, we present EgoHand, an proof-of-concept solution for VR interaction gesture recognition utilizing millimeter-wave radar and IMUs worn on the head for preserving more personal privacy.
-We conduct extensive cross-domain and few-shot experiments to evaluate EgoHand, highlighting the importance of training with samples from both hands to enhance cross-hand stability.
-Additionally, at least two pre-use calibrations are essential to ensure optimal performance across various users, scenes, and postures.
-Current EgoHand appears somewhat cumbersome, particularly with the TI mmWave radar.
-This can potentially be improved, for example, by using Google’s Soli radar (9mm×$\times$9mm) [23].
-In addition, mmWave radar may expose health metrics like breathing and heart rate [44, 4], which can be mitigated by adjusting the range and angle of radar signals.
+This paper presents mmEgoHand, a proof-of-concept system for egocentric hand pose estimation and gesture recognition, leveraging head-mounted millimeter-wave radar and IMUs.
+This configuration facilitates user mobility while offering enhanced personal privacy compared to traditional camera-based solutions.
+We conduct extensive experiments to evaluate mmEgoHand, demonstrating the importance of including samples from both hands during training to improve cross-hand generalization.
+Furthermore, we find that performing at least two pre-use calibration sessions is critical to achieving robust performance across diverse users, environments, and hand postures.
+While mmEgoHand  shows promising results, the current prototype is somewhat cumbersome, especially due to the use of the TI mmWave radar module.
+Future versions could benefit from more compact hardware, such as Google’s Soli radar (9mm×$\times$9mm) [50].
+Additionally, mmWave radar has the potential to reveal health-related metrics such as respiration and heart rate [51, 52], which raises privacy concerns.
+These can be mitigated by appropriately limiting the radar’s range and field of view.
+Moreover, although mmEgoHand  achieves promising accuracy in gesture recognition, there remains room for improvement in the precision of hand pose estimation and its generalization under cross-domain scenarios.
+To facilitate further research and development, we have publicly released our dataset, enabling the community to explore, benchmark, and improve upon our work.